A recipe says to add 2 cups of sugar per 5 cups of milk. if you already put in 5 cups of sugar, how much milk do you need to add?
Answer by mananth(16946)

let x be the number of cups required for 5 cups of sugar
the ratio of sugar to milk = 2/5
2/5 = 5/x
multiply by x
2/5 * x = 5x/x
2/5 x= 5
multiply by 5
2/5 *5 *x = 5*5
2x=25
divide by2
2x/2 = 25/2
x= 25/2
= 12.5 cups of milk required